Ina metre bridge experiment,when the resistances in the gaps are interchanged the balancing point increases by 10cm.the ratio of resistances will be ?

Grade:12th pass

1 Answers

Susmita
5 years ago
X/R=L/(100-L)
Now L changes to L'=L+10 when X and R are interchanged.Now
R/X=L'/(100-L')=(L+10)/(90-L)
Multiply the two equation.
1=L(L+10)/(100-L)(90-L)
Or,L2+10L=L2-190L+9000
Or,200L=9000
Or,L=45
So X/R=45/55=9/11
If it helps please approve.thank you.